400cm3 of helium gas effuses from a pourous container in 20 sec. How long will SO2 gas take to effuse from the same container?
1
Expert's answer
2020-12-21T03:53:26-0500

Time taken for Helium gas to effuse = 10 sec


Pressure, Volume, Temperature is constant since the same container is used.


Molar mass of Helium = 4 g/mol

Molar mass of SO2 = 64g/mol


According to graham's law of diffusion/effusion

The rate of diffusion of a gas is proportional to the square root of the inverse of the molar mass.


"r = \sqrt{\dfrac{1}{M}}"



"\dfrac{r_{He}}{r_{SO_2}} = \sqrt{\dfrac{M_{SO_2}}{M_{He}}}"


but, the rate is inversely proportional to time


"\dfrac{t_{SO_2}}{t_{He}} = \sqrt{\dfrac{M_{SO_2}}{M_{He}}}"


"\dfrac{t_{SO_2}}{10} = \sqrt{\dfrac{64}{4}}"


"\dfrac{t_{SO_2}}{10} = 4"


"t_{SO_2} = 40"


Therefore, it will take 40 seconds for "SO_2" gas take to effuse from the same container.